OpenStudy (anonymous):

8t^-3*t^-11 over t^-14

OpenStudy (anonymous):

Isn't this the same question? Answer is 8.

OpenStudy (anonymous):

t^3(8t^8 - 3)

OpenStudy (anonymous):

cant be 8?

jimthompson5910 (jim_thompson5910):

t^-3*t^-11 = t^(-3-11) = t^(-14) Divide this by t^(-14) in the denominator to get 1 So you then get 8*1 = 8

OpenStudy (anonymous):

If you really don't believe us sub any number as t into the equation (using a calculator) :)

jimthompson5910 (jim_thompson5910):

or you can graph it (replace t with x first)

jimthompson5910 (jim_thompson5910):

you'll get a horizontal line through y = 8 on the y-axis